OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal! Question- Simplify the following exponential expressions. (-.07xy^4)(.3x^3y)

OpenStudy (freckles):

multiply -.07 and .3 multiply x amd x^3 multply y^4 and y

OpenStudy (freckles):

use law of exponents for the last two

OpenStudy (daniel2raw):

then you will find your answer

OpenStudy (abbycross167):

Like this @freckles ? (-.07xy^4)(.3x^3y) = -.07*.3 = -0.021 x*x^3= 2x^3 y^4*y= 2y^y4

OpenStudy (freckles):

first product is correct the last two products you found is incorrect where is the 2 coming from in front if you have 1 factor of x and then you put another 3 factors of x with it how many factors of x do you have?

OpenStudy (abbycross167):

@jhonyy9 could you help me finish this? frecles is offline at the moment... Please?

jhonyy9 (jhonyy9):

x*x^3 = ? so than you know that a*a^2 = a^(1+2) =a^3

jhonyy9 (jhonyy9):

abby are you here ?

OpenStudy (abbycross167):

So x*x^3= x^4?

jhonyy9 (jhonyy9):

yes exactly you are right

jhonyy9 (jhonyy9):

and y^4 *y = ?

OpenStudy (abbycross167):

y*5?

jhonyy9 (jhonyy9):

yes so now you have got the right answer sure - yes ?

jhonyy9 (jhonyy9):

ok. ?

jhonyy9 (jhonyy9):

bye bye

OpenStudy (abbycross167):

I'm sorry my laptop messed up... is the answer -0.021x^4y^5?

jhonyy9 (jhonyy9):

yes exactly

OpenStudy (abbycross167):

thank you for the help!! :)

jhonyy9 (jhonyy9):

was my pleasure good luck bye bye